# Empirical Bayes for Data integration

### v1

@@ Abstract.
We discuss the use of empirical Bayes for data integration, in the sense of transfer learning. Our main interest is in settings where one wishes to learn structure (e.g. feature selection) and one only has access to incomplete data from previous studies, such as summaries, estimates or lists of relevant features. We discuss differences between full Bayes and empirical Bayes, and develop a computational framework for the latter. We discuss how empirical Bayes attains consistent variable selection under weaker conditions (sparsity and betamin assumptions) than full Bayes and other standard criteria do, and how it attains faster convergence rates. Our high-dimensional regression examples show that fully Bayesian inference enjoys excellent properties, and that data integration with empirical Bayes can offer moderate yet meaningful improvements in practice.
@@ 1. Introduction
There are two main scenarios where one may seek to integrate data, illustrated in Figure 1. The first (left panel) is when one has full access to multiple datasets. For example, one wishes to study multiple cancer types, and one has a separate dataset for each cancer type. Or, one has data from multiple schools and wants to learn common traits for low academic performance. We refer to this scenario as data integration with full data. There is a well-developed literature for this setting. A popular approach is to use hierarchical models that share information across datasets, see for instance Gelman and Hill (2007) for a monograph on mixed effects models. For data integration in covariance models see Avalos-Pacheco et al. (2022); Chandra et al. (2024), in clustering see Lock and Dunson (2013) and in functional data analysis see Yang et al. (2015). See also Jacob et al. (2017); Nott et al. (2023) and references therein for data integration strategies when the assumed model is misspecified.
Our main focus is in the second scenario (Figure 1, right), which occurs when one is interested in a single dataset $\bm{y}$, and one only has full access to that data. One also has other information that could be useful, for example summaries of previous datasets (denoted $\bm{Z}=(\bm{z}_{1},\ldots,\bm{z}_{q})$ in Figure 1). This second scenario is probably more common and certainly more challenging, because one does not have full access to the previous data and the provided summaries may be poor. An application that we illustrate later is a gene expression study in human colon cancer, where one has a list of genes that were found to be important in mice. The data analyst has full access to the human data, but only the list of genes from the mice experiment. We refer to this scenario as data integration with meta-covariates, i.e. with summaries $\bm{Z}$ that may be informative as to what parameters are more likely to play an important role in the main data of interest. Although this scenario received less attention than data integration with full data, there is a literature on the topic. To mention some examples, Stingo et al. (2011); Cassese et al. (2014) proposed Bayesian variable selection for gene expression where prior inclusion probabilities depend on biological knowledge and meta-covariates. Chen et al. (2021) predicted disease outcomes by allowing LASSO penalties to depend on functional annotations. In causal analysis, Belloni et al. (2012); Antonelli and Dominici (2021); Papaspiliopoulos et al. (2025) penalized the inclusion of control covariates using their degree of association with the treatment of interest. See van de Wiel et al. (2019) for an overview on the use of meta-covariates in regression. Beyond regression, in Gaussian graphical models Peterson et al. (2016) and Jewson et al. (2023) used node and edge covariates to drive prior edge inclusion probabilities. And in factor models, Schiavon et al. (2022) used meta-covariates to determine non-zero loadings.
@@
Figure 1. Data integration with full data (left) and with meta-covariates (right). With full data one observes $q$ datasets with underlying parameters $\bm{\theta}_{1},\ldots,\bm{\theta}_{q}$ with a hyper-parameter that allows sharing information. With meta-covariates one has access to summaries $\bm{z}_{1},\ldots,\bm{z}_{q}$ extracted from unobserved data $\tilde{\bm{y}}_{1},\ldots,\tilde{\bm{y}}_{q}$. These summaries inform the parameter $\bm{\theta}$ governing the dataset of interest $\bm{y}$
@@
Bayesian statistics provide a natural framework to integrate prior data such as meta-covariates into the analysis pipeline. In principle it suffices to set a prior on the parameters of interest ($\bm{\theta}$ in Figure 1, right panel) that depends on the meta-covariates $\bm{Z}$ and some hyper-parameters $\bm{\omega}$. There may be situations where an expert is willing to set such a prior subjectively. For example a biologist believes that if a gene is known to be involved in gastrointestinal cancer, then it has 0.9 prior probability of also being involved in colorectal cancer. Alas this situation is rather rare: experts do not have such a precise knowledge of $\bm{\omega}$, and it is cumbersome to specify such a prior when there are $q>1$ meta-covariates. This important because, if the prior grossly failed to reflect the true effect of $\bm{Z}$ (i.e., the expert provided a poor prior for $\bm{\omega}$), then one may obtain worse inference than if $\bm{Z}$ were not used at all.
@@
Our goal is to explain how empirical Bayes can be a useful strategy for data integration. By modeling the relation between the meta-covariates $\bm{Z}$ and the parameters $\bm{\theta}$ governing the observed data $\bm{y}$, one may learn $\bm{\omega}$ and hence integrate $\bm{Z}$ in a way that’s less sensitive to misspecification. We remark that empirical Bayes can be seen as a frequentist method where the model-fitting criteria are based on Bayesian thinking, and as such it can help reconcile long-standing debates between frequentist and Bayesian inference, see for example Efron (2024) for a recent discussion.
@@
The paper is structured as follows. Section 2 reviews basic notions related to empirical Bayes and to the frequentist validity of Bayesian methods, in the more common setting of data integration with full data. We also point to some literature discussing the benefits of empirical Bayes, which largely focus on parameter estimation and the full data setting. In Section 3 we outline an empirical Bayes framework for data integration with meta-covariates. For concreteness we focus on variable selection, but the setup extends directly to more general settings. Section 4 summarizes the scarce theory on structural learning with meta-covariates. In particular, we discuss how a simple form of meta-covariate (partitioning the parameters into blocks) leads to consistent model selection in settings where it is otherwise mathematically impossible (milder sparsity or betamin conditions), and that it improves consistency rates. By model selection (sometimes referred to support recovery) we mean identifying the subset of covariates truly associated to $\bm{y}$. We claim no methodological novelty for Sections 23. They are however not just a survey, rather they reflect the views of the authors on using empirical Bayes for data integration. Similarly, the main novelty in Section 4 is adapting results from Rognon-Vael et al. (2025) to a Bayesian setting. In Section 5 we present some new results that relate to computational algorithms to estimate the hyper-parameters $\hat{\bm{\omega}}$. A practical challenge is that evaluating the log marginal likelihood $p(\bm{y}\mid\bm{\omega})$ requires a sum over models (e.g. $2^{p}$ models in regression). We prove that its gradient can be evaluated in a linear cost in $p$ in fairly general settings, propose an Expectation-Maximization (EM, Dempster et al. (1977)) algorithm, for which we derive simplified gradients and hessians. In Section 6 we present simulated examples and an application to colon cancer data. These illustrate that fully Bayesian inference enjoys excellent properties and that its results are often fairly aligned with those from empirical Bayes. However in some settings, including our colon cancer example, there can be moderate but non-negligible differences between the two frameworks. Section 7 concludes.
@@ 2. Empirical Bayes
We start by discussing a purely frequentist interpretation of Bayesian inference. The interpretation holds exactly when the prior matches a certain data-generating truth, and asymptotically under certain settings. We discuss how this motivates the use of empirical Bayes as an attempt to improve the frequentist properties of Bayesian inference, and explain how fully Bayesian inference cannot attain such an improvement. We end by mentioning pitfalls of empirical Bayes and how to ameliorate them.
@@
To interpret Bayesian inference from a frequentist standpoint, it may be easiest to consider an example. Consider a setting where repeated experiments are conducted, with data from experiment $b$ generated as $\bm{y}^{(b)}\sim p(\bm{y}^{(b)}\mid\bm{\theta}^{(b)})$, and where the data-generating parameters $\bm{\theta}^{(b)}$ are drawn from some distribution $p(\bm{\theta})$. The usual frequentist setting envisions conducting repeated studies with the same exact parameter $\bm{\theta}^{(1)}=\bm{\theta}^{(2)}=\ldots$ (i.e. $p(\bm{\theta})$ is a point mass), whereas here we envision a frequentist setting where each experiment has its own data-generating $\bm{\theta}^{(b)}\sim p(\bm{\theta})$. In a Bayesian analysis, one also specifies a prior $\pi(\bm{\theta})$, which in general will not match $p(\bm{\theta})$. For example, a pharmaceutical company conducts repeated experiments over time to assess the efficacy of drugs $b=1,2,\ldots$. One observes the number of cured patients $y^{(b)}\sim\mbox{Beta}(n^{(b)},\theta^{(b)})$ and the parameter of interest is the curing probability $\theta^{(b)}$. Suppose that, unknown to the company, truly $\theta^{(b)}\sim\mbox{Beta}(2,11)$ independently across $b$ (i.e., $p(\bm{\theta})$ is given by this Beta distribution). Figure 2 shows an illustration for $n=30$ and $n=100$. For simplicity it assumes that $n^{(1)}=n^{(2)}=\ldots=n$, and that the experiments are independent (given $\theta^{b}$). The left panels show the $\mbox{Beta}(2,11)$ density, two realizations $\theta^{(b)}=0.1$ and $0.35$ from this $p(\theta)$, and the corresponding Binomial probability mass functions. Suppose that the data analyst sets independent $\mbox{Beta}(a,b)$ priors to each $\theta^{(b)}$, denoted $\pi(\theta^{(b)})$ earlier, for some $a,b>0$. The right panels show the posterior distributions $\pi(\theta^{(b)}\mid y^{(b)})$ for two putative $y^{(b)}$ if the data analyst were to use the correct $\mbox{Beta}(2,11)$ prior (black solid lines), where by correct we mean matching the data-generating truth. These posteriors are the frequentist distribution of the $\theta^{(b)}$’s conditional on the observed $y^{(b)}$, that is $\pi(\theta^{(b)}\mid y^{(b)})=p(\theta^{(b)}\mid y^{(b)})$ where $p(\theta^{(b)}\mid y^{(b)})\propto p(y^{(b)}\mid\theta^{(b)})p(\theta^{(b)})$. No Bayesian interpretation is needed: $(\theta^{(b)},y^{(b)})$ have a frequentist distribution, $p(\theta^{(b)}\mid y^{(b)})$ is a well-defined frequentist object, and posterior inference is well-calibrated from a frequentist standpoint.
@@
Figure 2 also shows another posterior that only has a Bayesian interpretation, obtained by setting a $\mbox{Beta}(4,9)$ prior (dashed grey lines). The new posterior no longer has a direct frequentist interpretation under the data-generating $p(\bm{\theta})$, but as $n$ grows (bottom panels) both posteriors converge to the same answer by virtue of the Bernstein-von Mises theorem (see van der Vaart (1998), Theorem 10.1, for example). That is, even if one sets the wrong (in a frequentist sense) prior, when certain limiting results apply the posterior asymptotically has a valid frequentist interpretation. Alas, such limiting theorems typically apply to regular finite-dimensional models, and high-dimensional settings where $\mbox{dim}(\bm{\theta}^{(b)})\gg n$ must be handled with more care.
@@ 3. Framework
We lay out an empirical Bayes framework for data integration via meta-covariates. The framework is not new, see van de Wiel et al. (2019) and references therein. Consider a setting where the main data records an outcome $\bm{y}=(y_{1},\ldots,y_{n})^{T}\in\mathbb{R}^{n}$ for $n$ individuals and $p$ covariates $\bm{x}_{i}\in\mathbb{R}^{p}$ for $i=1,\ldots,n$. For each covariate $j=1,\ldots,p$, one also has a vector of $q$ meta-covariates $\bm{z}_{j}\in\mathbb{R}^{q}$. We denote by $\bm{X}$ the $n\times p$ covariates matrix with $i^{th}$ row equal to $\bm{x}_{i}$, and by $\bm{Z}$ the $p\times q$ metacovariates matrix with $\bm{z}_{j}$ as its $j^{th}$ row. We remark that the framework can be easily extended to other settings such as graphical models or factor models (Peterson et al., 2016; Schiavon et al., 2022; Jewson et al., 2023), but for simplicity here we focus on regression.
Before proceeding, we discuss some examples to make ideas concrete. Suppose that one wants to study the association between colon cancer survival ($y_{i}$) and gene expression ($\bm{x}_{i}$), and one has three lists of genes found to be associated to gastrointestinal, endometrial and ovarian cancer survival. Here, $\bm{z}_{j}\in\{0,1\}^{3}$ indicates whether gene $j$ is included in these 3 lists. The meta-covariates $\bm{z}_{j}$ may contain further information, such as the biological pathways that each gene is known to be associated with, or whether the covariate corresponds to clinical variable, known risk factors, genomic data, etc. As another example, the director of a school wishes to identify socio-demographic covariates $\bm{X}$ associated with academic performance $\bm{y}$. The director has reports giving estimated covariate effects $\bm{Z}$ in other schools, but does not have access to the full data from those schools.
@@
As discussed, the key idea is to set a prior $\pi(\bm{\theta}\mid\bm{Z},\bm{\omega})$ that depends on the meta-covariates $\bm{Z}$ via hyper-parameters $\bm{\omega}$, and where $\bm{\omega}$ is estimated via empirical Bayes. For simplicity we denote $\pi(\bm{\theta}\mid\bm{\omega}):=\pi(\bm{\theta}\mid\bm{Z},\bm{\omega})$. We are primarily interested in structural learning where one considers (possibly many) different models for $\bm{\theta}$. We denote such models generically by $\bm{\gamma}\in\Gamma$, where $\Gamma$ is the set of possible models. For example, in variable selection we have $\bm{\theta}=(\theta_{1},\ldots,\theta_{p})^{T}$ regression parameters and the models are indexed by $\bm{\gamma}=(\gamma_{1},\ldots,\gamma_{p})\in\{0,1\}^{p}$, where $\gamma_{j}=\mbox{I}(\theta_{j}\neq 0)$ is an inclusion indicator for covariate $j=1,\ldots,p$. One may also have parameters that are present in all models, such as the error variance or covariates that one forces to be included in the model, but we omit these from the notation for simplicity. Formally, each model $\bm{\gamma}$ has its own likelihood $p(\bm{y}\mid\bm{\theta},\bm{\gamma})$, and the joint fully Bayesian model is
$p(\bm{y},\bm{\theta},\bm{\gamma})=p(\bm{y}\mid\bm{\theta},\bm{\gamma})\pi(\bm{\theta}\mid\bm{\gamma})\pi(\bm{\gamma}),$
where $\pi(\bm{\gamma})$ is some given model prior. Using Bayes formula, posterior model probabilities are given by $p(\bm{\gamma}\mid\bm{y})\propto p(\bm{y}\mid\bm{\gamma})\pi(\bm{\gamma})$, where
@@
is the so-called marginal likelihood of model $\bm{\gamma}$, and $\Pi(\bm{\theta}\mid\bm{\gamma})$ the prior measure on the parameters under $\bm{\gamma}$.
To achieve data integration within full Bayes, one would replace $\pi(\bm{\gamma})$ by some $\pi(\bm{\gamma}\mid\bm{\omega})$ that depends on the meta-covariates $\bm{Z}$, and then set a hyper-prior $\pi(\bm{\omega})$. For example, one may set prior inclusion probabilities $\pi(\gamma_{j}=1\mid\bm{\omega})$ that depend on the $j^{th}$ parameter’s covariates $\bm{z}_{j}$ via a logistic regression with linear predictor $\bm{z}_{j}^{T}\bm{\omega}$. The issue is that such a fully Bayesian framework gives posterior probabilities $\pi(\bm{\gamma}\mid\bm{y})\propto p(\bm{y}\mid\bm{\gamma})\pi(\bm{\gamma})$, where
@@
By definition $\pi(\bm{\gamma})$ does not depend on $\bm{y}$, it is a quantity defined a priori, and hence it cannot take advantage of learning $\hat{\bm{\omega}}$ from $\bm{y}$. That is, although such $\pi(\bm{\gamma})$ depends on meta-covariates, this dependence is set entirely a priori. For example, suppose that one wants to set different prior inclusion probabilities $\pi(\gamma_{j}=1)$ for each $j$, but one wishes to learn them from the data, and sets an exchangeable $\pi(\bm{\gamma}\mid\bm{\omega})$. This task it not possible because $\pi(\bm{\gamma})$ is exchangeable (it’s a mixture of exchangeable $\pi(\bm{\gamma}\mid\bm{\omega})$).

### v2

@@ Abstract.
We discuss the use of empirical Bayes for data integration, in the sense of transfer learning. Our main interest is in settings where one wishes to learn structure (e.g., feature selection) and one only has access to incomplete data from previous studies, such as summaries, estimates or lists of relevant features. We discuss differences between full Bayes and empirical Bayes, and develop a computational framework for the latter. We discuss how empirical Bayes attains consistent variable selection under weaker conditions (sparsity and betamin assumptions) than full Bayes and other standard criteria do, and how it attains faster convergence rates. Our high-dimensional regression examples show that fully Bayesian inference enjoys excellent properties, and that data integration with empirical Bayes can offer moderate yet meaningful improvements in practice.
@@ 1. Introduction
There are two main scenarios where one may seek to integrate data, illustrated in Figure 1. The first (left panel) is when one has full access to multiple datasets. For example, one has a separate dataset for multiple cancer types, or for multiple schools. We refer to this scenario as data integration with full data. There is a well-developed literature for this setting. A popular approach is to use hierarchical models that share information across datasets, see for instance Gelman and Hill (2007) for a monograph. For data integration in covariance models see Avalos-Pacheco et al. (2022); Chandra et al. (2024), in clustering see Lock and Dunson (2013) and in functional data analysis see Yang et al. (2015). See also Jacob et al. (2017); Nott et al. (2023) and references therein for data integration strategies when the assumed model is misspecified.
Our main focus is in the second scenario (Figure 1, right), where one is interested in a single dataset $\bm{y}$, and one only has full access to that data. One also has other information that could be useful, for example summaries of previous datasets (denoted $\bm{Z}=(\bm{z}_{1},\ldots,\bm{z}_{q})$ in Figure 1). This second scenario is very common and more challenging, because one does not have full access to the previous data and the provided summaries may be poor. An application that we illustrate later is a gene expression study in human colon cancer, where one has a list of genes that were found to be important in mice. We have full access to the human data, but only the list of genes from the mice experiment. We refer to this scenario as data integration with meta-covariates, i.e. with summaries $\bm{Z}$ that may be informative as to what parameters are more likely to be important. Literature focusing on this scenario includes Stingo et al. (2011); Cassese et al. (2014), who proposed Bayesian variable selection where prior inclusion probabilities depend on biological knowledge and meta-covariates. Chen et al. (2021) predicted disease outcomes by allowing LASSO penalties to depend on functional annotations. In causal analysis, Belloni et al. (2012); Antonelli and Dominici (2021); Papaspiliopoulos et al. (2025) penalized the inclusion of control covariates using their degree of association with the treatment of interest. See van de Wiel et al. (2019) for an overview on the use of meta-covariates in regression. Beyond regression, in Gaussian graphical models Peterson et al. (2016) and Jewson et al. (2023) used node and edge covariates to drive prior edge inclusion probabilities. In factor models, Schiavon et al. (2022) used meta-covariates to determine non-zero loadings.
@@
Figure 1. Data integration with full data (left) and with meta-covariates (right). With full data one observes $q$ datasets with underlying parameters $\bm{\theta}_{1},\ldots,\bm{\theta}_{q}$. The hyper-parameter $\bm{\omega}$ allows sharing information. With meta-covariates one has access to summaries $\bm{z}_{1},\ldots,\bm{z}_{q}$ extracted from unobserved data $\tilde{\bm{y}}_{1},\ldots,\tilde{\bm{y}}_{q}$. These summaries inform the parameter $\bm{\theta}$ governing the dataset of interest $\bm{y}$
@@
Bayesian statistics provides a natural framework to integrate prior data such as meta-covariates $\bm{Z}$ into the analysis. In principle, it suffices to set a prior on the parameters of interest ($\bm{\theta}$ in Figure 1, right panel) that depends on $\bm{Z}$ and some hyper-parameters $\bm{\omega}$. There may be situations where an expert is willing to set such a prior subjectively. For example a biologist believes that if a gene is known to be involved in gastrointestinal cancer, then it has 0.9 prior probability of also being involved in colorectal cancer. This situation is rare. Experts do not have such a precise knowledge, specially when there are $q>1$ meta-covariates. This is important: if the prior grossly failed to reflect the true effect of $\bm{Z}$ (i.e., the expert provided a poor prior for $\bm{\omega}$), then one may obtain worse inference than if $\bm{Z}$ were not used at all.
@@
We explain how empirical Bayes can be useful for data integration. By modeling the relation between the meta-covariates $\bm{Z}$ and the parameters $\bm{\theta}$ governing the data $\bm{y}$, one may learn $\bm{\omega}$ and hence integrate $\bm{Z}$ in a way that’s less sensitive to misspecification. Empirical Bayes can be seen as a frequentist method where the model-fitting criteria are based on Bayesian thinking, and as such it can help reconcile long-standing debates between frequentist and Bayesian inference, see for example Efron (2024).
@@
The paper is structured as follows. Section 2 reviews empirical Bayes and the frequentist validity of Bayesian methods, in the setting of data integration with full data. In Section 3 we outline an empirical Bayes framework for data integration with meta-covariates. For concreteness, we focus on variable selection, but the setup extends directly to more general settings. Section 4 provides theory on high-dimensional linear regression with meta-covariates. We discuss how a simple form of meta-covariate (partitioning covariates into blocks) leads to consistent model selection (identifying the truly non-zero parameters) in settings where it is otherwise mathematically impossible (milder sparsity or betamin conditions), and that it improves consistency rates. Sections 2-3 are not methodologically new, but they reflect the views of the authors on using empirical Bayes for data integration. The main novelty in Section 4 is adapting results from Rognon-Vael et al. (2025) to a Bayesian setting. Section 5 includes new computational results to estimate the hyper-parameters $\hat{\bm{\omega}}$ that apply to very general settings, and in particular beyond regression. Evaluating the log marginal likelihood $p(\bm{y}\mid\bm{\omega})$ requires a sum over models (e.g. $2^{p}$ in regression), but its gradient can be expressed as a sum over $p$ terms, and we propose an Expectation-Maximization (EM, Dempster et al. (1977)) algorithm. In Section 6 we present simulations and an application to colon cancer data. These illustrate that fully Bayesian inference enjoys excellent properties and that its results are often aligned with those from empirical Bayes. However in some settings, including our colon cancer example, there can be moderate but non-negligible differences between the two frameworks. Section 7 concludes.
@@ 2. Empirical Bayes
We start by discussing a purely frequentist interpretation of Bayesian inference. The interpretation holds exactly when the prior matches a certain data-generating truth, and asymptotically under certain settings. We discuss how this motivates using empirical Bayes to try to improve the frequentist calibration of fully Bayesian inference. We end by mentioning pitfalls of empirical Bayes and how to ameliorate them.
@@
Interpreting Bayesian inference from a frequentist standpoint may be easiest with an example. Consider a setting where repeated experiments are conducted, with data from experiment $b$ generated as $\bm{y}^{(b)}\sim p(\bm{y}^{(b)}\mid\bm{\theta}^{(b)})$, and where the data-generating parameters $\bm{\theta}^{(b)}$ are drawn from some distribution $p(\bm{\theta})$. The usual frequentist setting envisions repeated studies with the same parameter $\bm{\theta}^{(1)}=\bm{\theta}^{(2)}=\ldots$, whereas we envision a frequentist setting where each experiment has its own data-generating $\bm{\theta}^{(b)}\sim p(\bm{\theta})$. In a Bayesian analysis, one also specifies a prior $\pi(\bm{\theta})$, which in general will not match $p(\bm{\theta})$. For example, a pharmaceutical company conducts repeated experiments over time to assess the efficacy of drugs $b=1,2,\ldots$. One observes the number of cured patients $y^{(b)}\sim\mbox{Beta}(n^{(b)},\theta^{(b)})$ and the parameter of interest is the cure probability $\theta^{(b)}$. Suppose that, unknown to the company, truly $\theta^{(b)}\sim\mbox{Beta}(2,11)$ independently across $b$, denoted by $p(\theta)$ earlier. Figure 2 shows an illustration for $n=30$ and $n=100$. For simplicity, it assumes that $n^{(1)}=n^{(2)}=\ldots=n$, and that the experiments are independent given $\theta^{(b)}$. The left panels show the $\mbox{Beta}(2,11)$ density, two realizations $\theta^{(b)}=0.1$ and $0.35$, and the corresponding Binomial probability mass functions. Suppose that one sets independent $\mbox{Beta}(a,l)$ priors to each $\theta^{(b)}$, denoted $\pi(\theta^{(b)})$ earlier, for some $a,l>0$. The right panels show the posterior distributions $\pi(\theta^{(b)}\mid y^{(b)})$ for two putative $y^{(b)}$ if the data analyst were to use the $\mbox{Beta}(2,11)$ prior (black solid lines) matching the data-generating truth. These posteriors are the frequentist distribution of the $\theta^{(b)}$’s conditional on the observed $y^{(b)}$, that is $\pi(\theta^{(b)}\mid y^{(b)})=p(\theta^{(b)}\mid y^{(b)})$ where $p(\theta^{(b)}\mid y^{(b)})\propto p(y^{(b)}\mid\theta^{(b)})p(\theta^{(b)})$. No Bayesian interpretation is needed: $(\theta^{(b)},y^{(b)})$ have a frequentist distribution, $p(\theta^{(b)}\mid y^{(b)})$ is a well-defined frequentist object, and posterior inference is well-calibrated from a frequentist standpoint.
@@
Figure 2 also shows another posterior that only has a Bayesian interpretation, obtained by setting the prior $\pi(\theta^{(b)})$ to a $\mbox{Beta}(4,9)$ (dashed grey lines). The new posterior no longer has a direct frequentist interpretation under the data-generating $p(\bm{\theta})$, but as $n$ grows (bottom panels) both posteriors converge to the same answer by virtue of the Bernstein-von Mises theorem (van der Vaart (1998), Theorem 10.1). That is, even if one sets the wrong prior $\pi(\theta^{(b)})$ (in a frequentist sense, that is not matching $p(\theta^{(b)})$), when certain limiting results apply the posterior asymptotically has a valid frequentist interpretation. However, such limiting theorems typically apply to finite-dimensional models, and high-dimensional settings must be handled with care.
@@ 3. Framework
We lay out an empirical Bayes framework for data integration via meta-covariates, akin to van de Wiel et al. (2019). Consider a setting where the main data records an outcome $\bm{y}=(y_{1},\ldots,y_{n})^{T}\in\mathbb{R}^{n}$ for $n$ individuals and $p$ covariates $\bm{x}_{i}\in\mathbb{R}^{p}$ for $i=1,\ldots,n$. For each covariate $j=1,\ldots,p$, one also has a vector of $q$ meta-covariates $\bm{z}_{j}\in\mathbb{R}^{q}$. We denote by $\bm{X}$ the $n\times p$ covariates matrix with $i^{th}$ row equal to $\bm{x}_{i}$, and by $\bm{Z}$ the $p\times q$ meta-covariates matrix with $\bm{z}_{j}$ as its $j^{th}$ row. The framework can be extended to other settings such as covariance models (Jewson et al., 2023), but for simplicity we focus on regression.
Before proceeding, we discuss examples to make ideas concrete. Suppose that one wants to study the association between colon cancer survival ($y_{i}$) and gene expression ($\bm{x}_{i}$), and one has three lists of genes found to be associated with gastrointestinal, endometrial and ovarian cancer. Here, $\bm{z}_{j}\in\{0,1\}^{3}$ indicates whether gene $j$ is included in these 3 lists. The meta-covariates $\bm{z}_{j}$ may contain further information: biological pathways in which each gene features, whether the covariate is a clinical or genomic variable, etc. As another example, a school’s director wants to identify covariates $\bm{X}$ associated with academic performance $\bm{y}$. The director has reports $\bm{Z}$ on estimated covariate effects in other schools, but does not have the full data from those schools.
@@
The key idea is to set a prior $\pi(\bm{\theta}\mid\bm{Z},\bm{\omega})$ that depends on the meta-covariates $\bm{Z}$ via hyper-parameters $\bm{\omega}$, and where $\bm{\omega}$ is estimated via empirical Bayes. For simplicity, we denote $\pi(\bm{\theta}\mid\bm{\omega}):=\pi(\bm{\theta}\mid\bm{Z},\bm{\omega})$. We are interested in structural learning where one considers many different models for $\bm{\theta}$. We denote such models generically by $\bm{\gamma}\in\Gamma$, where $\Gamma$ is the set of models. For example, in variable selection we have $\bm{\theta}=(\theta_{1},\ldots,\theta_{p})^{T}$ regression parameters and the models are indexed by $\bm{\gamma}=(\gamma_{1},\ldots,\gamma_{p})\in\{0,1\}^{p}$, where $\gamma_{j}=\mbox{I}(\theta_{j}\neq 0)$ is an inclusion indicator for covariate $j=1,\ldots,p$. One may also have parameters that are present in all models, such as the error variance or covariates not undergoing selection, but we omit these from the notation for simplicity.
Formally, each model $\bm{\gamma}$ has its own likelihood $p(\bm{y}\mid\bm{\theta},\bm{\gamma})$, and the joint Bayesian model is $p(\bm{y},\bm{\theta},\bm{\gamma})=p(\bm{y}\mid\bm{\theta},\bm{\gamma})\pi(\bm{\theta}\mid\bm{\gamma})\pi(\bm{\gamma}),$ where $\pi(\bm{\gamma})$ is a given model prior. Using Bayes formula, posterior model probabilities are $p(\bm{\gamma}\mid\bm{y})\propto p(\bm{y}\mid\bm{\gamma})\pi(\bm{\gamma})$, where
@@
is the marginal likelihood of model $\bm{\gamma}$, and $\Pi(\bm{\theta}\mid\bm{\gamma})$ the prior measure on $\bm{\theta}$ under $\bm{\gamma}$.
To achieve data integration within full Bayes, one would replace $\pi(\bm{\gamma})$ by some $\pi(\bm{\gamma}\mid\bm{\omega})$ that depends on the meta-covariates $\bm{Z}$, and then set a hyper-prior $\pi(\bm{\omega})$. Although much of our discussion and some of our results are more general, to fix ideas consider $\pi(\bm{\gamma}\mid\bm{\omega})=\prod_{j=1}^{p}\mbox{Bern}(\gamma_{j};m_{j}(\bm{\omega}))$ for a given differentiable $m_{j}(\bm{\omega})$, such as the inverse logit function $m_{j}(\bm{\omega})=1/(1+e^{-\bm{z}_{j}^{T}\bm{\omega}})$. In what follows, we indicate where this prior structure is explicitly assumed. The issue is that such a fully Bayesian framework gives posterior probabilities $\pi(\bm{\gamma}\mid\bm{y})\propto p(\bm{y}\mid\bm{\gamma})\pi(\bm{\gamma})$, where
@@
By definition $\pi(\bm{\gamma})$ does not depend on $\bm{y}$, it is a quantity defined a priori, and hence by definition it is not data-adaptive. That is, although $\pi(\bm{\gamma})$ depends on meta-covariates, this dependence is set entirely a priori. This statement may sound surprising, given the popular practice of placing priors on hyper-parameters such as $\bm{\omega}$, with the hope of learning $\bm{\omega}$ and guiding posterior inference via $\pi(\bm{\omega}\mid\bm{y})$. Specifically, since
(1) $\pi(\bm{\gamma}\mid\bm{y})=\int\pi(\bm{\gamma}\mid\bm{y},\bm{\omega})\pi(\bm{\omega}\mid\bm{y})d\bm{\omega}$
where $\pi(\bm{\omega}\mid\bm{y})$ depends on $\bm{y}$, it should follow that $\pi(\bm{\gamma}\mid\bm{y})$ adapts to learning $\bm{\omega}$ from $\bm{y}$. This interpretation of (1) is misleading. As noted $\pi(\bm{\gamma}\mid\bm{y})\propto p(\bm{y}\mid\bm{\gamma})\pi(\bm{\gamma})$ where $\pi(\bm{\gamma})$ does not depend on $\bm{y}$, hence $\pi(\bm{\gamma}\mid\bm{y})$ can only adapt to $\bm{y}$ via $p(\bm{y}\mid\bm{\gamma})$. To reconcile this observation with (1), using that $\pi(\bm{\gamma}\mid\bm{y},\bm{\omega})=p(\bm{y}\mid\bm{\gamma},\bm{\omega})\pi(\bm{\gamma}\mid\bm{\omega})/p(\bm{y}\mid\bm{\omega})$ and that $\pi(\bm{\omega}\mid\bm{y})=p(\bm{y}\mid\bm{\omega})\pi(\bm{\omega})/p(\bm{y})$, re-write (1) as $\pi(\bm{\gamma}\mid\bm{y})=$
(2) $\int\frac{p(\bm{y}\mid\bm{\gamma},\bm{\omega})\pi(\bm{\gamma}\mid\bm{\omega})}{p(\bm{y}\mid\bm{\omega})}\frac{p(\bm{y}\mid\bm{\omega})\pi(\bm{\omega})}{p(\bm{y})}d\bm{\omega}=\frac{\int p(\bm{y}\mid\bm{\gamma},\bm{\omega})\pi(\bm{\gamma}\mid\bm{\omega})\pi(\bm{\omega})d\bm{\omega}}{p(\bm{y})}=\frac{p(\bm{y}\mid\bm{\gamma})\pi(\bm{\gamma})}{p(\bm{y})}$
since $p(\bm{y}\mid\bm{\gamma},\bm{\omega})=p(\bm{y}\mid\bm{\gamma})$. Critically, in the left-hand side of (2) the term $p(\bm{y}\mid\bm{\omega})$ coming from $\pi(\bm{\omega}\mid\bm{y})$, which is how one learns hyper-parameters from $\bm{y}$, cancels with $p(\bm{y}\mid\bm{\omega})$ coming from $\pi(\bm{\gamma}\mid\bm{\omega},\bm{y})$. That is, is true that $\pi(\bm{\omega}\mid\bm{y})$ learns about $\bm{\omega}$ from $\bm{y}$, but the effect of such learning is cancelled in $\pi(\bm{\gamma}\mid\bm{y})$ because $p(\bm{y}\mid\bm{\omega})$ also features in $\pi(\bm{\gamma}\mid\bm{\omega},\bm{y})$. These arguments are not restricted to Bayesian model selection, they apply to any hierarchical model where one sets a prior on hyper-parameters.